OpenStudy (anonymous):

Use the elimination system of linear equation.(my three key is broke..) a+7b=1 a+ three b= negative three

OpenStudy (anonymous):

This is like the two we already did isnt it?

OpenStudy (anonymous):

yeah bu this has mor letters like the other ones had letters with the number

OpenStudy (anonymous):

Ok you wanna eliminate one of the variables so lets multiply the first equation by -1. so now we have: -a-7b=-1 a+3b=-3 now add the equations together and solve for b.

OpenStudy (anonymous):

ok so i add 7b to negative three

OpenStudy (anonymous):

because i believe -a and a is canceled out

OpenStudy (anonymous):

No you add the like terms. straight down. youll get: -4b=-4

OpenStudy (anonymous):

oh ok 7 minus negativ three and 1+ three ok

OpenStudy (anonymous):

b=1

OpenStudy (anonymous):

i divided

OpenStudy (anonymous):

Yep. Now that you know b, plug it in to one of the original equations and solve for a

OpenStudy (anonymous):

so, i place 1 for a

OpenStudy (anonymous):

No you put in 1 for b.

OpenStudy (anonymous):

a+7(1)=1 solve for a

OpenStudy (anonymous):

it seems to be three and 7 for th solution becausee the letter a in this equation is 1

OpenStudy (anonymous):

correction (1,7)

OpenStudy (anonymous):

You should get -6 for a.

OpenStudy (anonymous):

7(1)=7 so minus 1= 6

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

(1,-6)